OpenStudy (anonymous):

Can someone help me? This problem is worth 20 points towards my next exam :( A bus starts from rest and accelerates at 2 m/sec2 until it reaches a velocity of 12 m/sec. The bus continues at this velocity for a period of time and then decelerates at 2.5 m/sec2 and comes to a stop 900 m. from it’s starting point. How much time did the bus take to travel the 900 m?

OpenStudy (stormfire1):

First part of trip: a = 2 m/s^2 Vi = 0 m/s Vf = 12 m/s Xi = 0 Xf = ? t = ? Now, find the unknowns. Since it's constant acceleration, to calculate the time it took to get to 12m/s use the formula: Vf = Vi + a*t 12 m/s = 0 m/s + (2 m/s^2) * t t = 6 seconds Now see how far it went in those 6 seconds: Xf - Xi = Vi * t + .5a * t^2 Xf - 0m = 0 m/s + .5(2m/s^2)(6s^2) Xf = 36m 900m - 36m = 864m to go... Now you need to figure out how long it will take it to slow down if it's going 12m/s with a deceleration of -2.5m/s^2. a = -2.5 m/s^2 Vi = 12 m/s Vf = 0 m/s Xi = 0 Xf = ? t = ? Similar to before: Vf = Vi + a * t 0m/s = 12m/s + (-2.5m/s^2)*t t = 4.8s Now see how far it took it to stop: Xf - Xi = Vi * t + .5a * t^2 Xf - 0m = 12m/s * 4.8s + .5(-2.5m/s^2)(4.8s^2) Xf = 28.8m So, so how much distance is left? 900m - 36m - 28.8m = 835.2m It was going 12 m/s for the entire time it took it to go 835.2m...so it took it: (835.2m)/(12m/s) = 69.6 seconds Add up the times to get the final answer: 69.6s + 6s + 4.8s = 80.4s Yep, I can see why this one was extra credit :P

OpenStudy (anonymous):

Thank you so much! I finally get it now :) my professor made us buy a cheap 20 dollar physics book that has all the examples but doesn't show the work and how they got the answer :( thanks again, you're a life saver!

OpenStudy (stormfire1):

Glad to help :)